# On the Multiary Algebraic Formulation of an Idempotent Symmetric Limit Convex Structure

### v1

@@ 4.2 Separation of $\mathbb{B}$-Convex Sets over $\mathbb{R}^{n}$
where $d_{H,\infty}$ denotes the Hausdorff distance induced by the $\|\cdot\|_{\infty}$ norm. One can then construct sequence $\{A_{m}\}_{m\in\mathbb{N}}$ a sequence of finite subsets of points of $C$, such that $A_{m}\subset A_{m+1}$ and $d_{H,\infty}(A_{m},C)\leq\frac{1}{m}$ for any positive integer $m$. Since $C$ is $\mathbb{B}$-convex, for any $m$, $Co^{\infty}(A_{m})$ is a subset of $C$. Therefore, since $C\supset Co^{\infty}(A_{m})\supset A_{m}$, we deduce that:
@@
In [11], it was claimed that an external representation of a $\mathbb{B}$-polytope can be given from a finite collection of lower halfspaces. In the following, we slightly correct this result, which only holds in the case where the intersection set is topologically regular.
@@
Proof: From Lemma [11], there exists an increasing subsequence $\{p_{q}\}_{q\in\mathbb{N}}$ and $\ell$ subsequences of linear forms $\{f_{j,p_{q}}\}_{q\in\mathbb{N}}$ respectively defined for all natural numbers $q$ by $f_{j,p_{q}}(x)=\langle a_{j}^{(p_{q})},x\rangle_{p_{q}}$ with ${{Co}}^{p_{q}}(A)=\bigcap_{j\in[\ell]}{\big[f_{j,p_{q}}\leq c_{j,p_{q}}\big]}$. Hence for each $j$:

### v2

@@ 4.2 Separation of $\mathbb{B}$-Convex Sets over $\mathbb{R}^{n}$
where $d_{H,\infty}$ denotes the Hausdorff-Pompeiu distance induced by the $\|\cdot\|_{\infty}$ norm. One can then construct sequence $\{A_{m}\}_{m\in\mathbb{N}}$ a sequence of finite subsets of points of $C$, such that $A_{m}\subset A_{m+1}$ and $d_{H,\infty}(A_{m},C)\leq\frac{1}{m}$ for any positive integer $m$. Since $C$ is $\mathbb{B}$-convex, for any $m$, $Co^{\infty}(A_{m})$ is a subset of $C$. Therefore, since $C\supset Co^{\infty}(A_{m})\supset A_{m}$, we deduce that:
@@
In [11], it was claimed that an external representation of a $\mathbb{B}$-polytope can be given from a finite collection of lower halfspaces. In the following, we slightly correct this result, which only holds in the case where the intersection set is topologically regular. A closed subset $E\subset\mathbb{R}^{n}$ is said to be topologically regular if $E=\mathrm{cl}\big({\operatorname{int}(E)}\big).$ In other words, $E$ coincides with the closure of its interior.
@@
Proof: From [11], there exists an increasing subsequence $\{p_{q}\}_{q\in\mathbb{N}}$ and $\ell$ subsequences of linear forms $\{f_{j,p_{q}}\}_{q\in\mathbb{N}}$ respectively defined for all natural numbers $q$ by $f_{j,p_{q}}(x)=\langle a_{j}^{(p_{q})},x\rangle_{p_{q}}$ with ${{Co}}^{p_{q}}(A)=\bigcap_{j\in[\ell]}{\big[f_{j,p_{q}}\leq c_{j,p_{q}}\big]}$. Hence for each $j$: